Consider a recurrence relation an = -3an-1 + an-2 for n = 1,2,3,4,… with initial conditions a= 4 and a= 2. Calculate a5.


1
Expert's answer
2021-11-24T17:29:42-0500

Consider a recurrence relation "a_n = -3a_{n-1} + a_{n-2}"   with initial conditions "a_1 = 4" and "a_2 = 2." Let us calculate "a_5." It follows that "a_3 = -3a_{2} + a_{1}=-3\cdot2+4=-2." Then "a_4 = -3a_{3} + a_{2}=-3\cdot(-2)+2=8." We conclude that "a_5 = -3a_{4} + a_{3}=-3\cdot 8-2=-26."

Answer: "a_5=-26."